OpenStudy (anonymous):

A rope is swinging in such a way that the length of the arc is decreasing geometrically. If the first arc is 18 feet long and the third arc is 8 feet long, what is the length of the second arc?

OpenStudy (anonymous):

are u dizzyy

OpenStudy (anonymous):

no

OpenStudy (anonymous):

let second rope be x..then x/18 = 8/x or x^2 = 18 x 8 = 144 x =12...this is the answer

OpenStudy (phi):

geometric means you multiply by some factor to get the next number you start at 18 the next number is 18*K (K is the factor) the 3rd number is 18*k*k but we know the third number is 8

OpenStudy (anonymous):

okay so would 12 be my answer ?

OpenStudy (anonymous):

yes ...for sure...12 is the answer

OpenStudy (phi):

you can figure it out 18*k*k = 8 divide by 18 on both sides k*k= 8/18 or 4/9 k= sqrt(4/9) = 2/3 the middle number is k*18

OpenStudy (anonymous):

I got it @phi Thank you & Thank you @ritez